Does shorter player shoot more free throws?

Calculate the 95% Confidence Interval(Interval Estimation) for the data below.?

    Player                                       Height                                Free Throw %

Giannis Antetokounmpo

                 6’11

                 72.9

Thanasis Antetokounmpo

                 6’6

                 50.0

Dragan Bender

                 7’0

                 65.0

Eric Bledsoe

                 6’1

                 79.4

Sterling Brown

                 6’5

                 74.7

Pat Connaughton

                 6’5

                 79.1

Donte DiVencenzo

                 6’4

                 76.3

George Hill

                 6’3

                 80.2

Ersan llyasova

                 6’9

                 77.4

Kyle Korver

                 6’7

                 87.7

Brook Lopez

                 7’0

                 79.3

Robert Lopez

                 7’0

                 75.5

Frank Mason III

                 5’11

                 77.0

Wesley Matthews

                 6’4

                 82.6

Kris Middleton

                 6’7

                 87.0

Cameron Reynolds

                 6’7

                 88.9

D.J. Wlison

                 6’10

                 56.0
Answer #1

In this problem first I have converted the given height in inches as follows-

we know , 1 foot =12 inches . Thus, 6'11=(6*12)+11=72+11=83 inch. Similarly for others.U Here, we have to test whether shorter player shoot more free throws? se we have one sided alternative H, le call Where l -Now to obtain X bar , Ybar and pooled variance I had used EXCEL as follows=

Data - Data analysis - 2 sample t test. It gives output as follws:

t-Test Two-Sample Assuming Equal Variances height(in inches) free throw 83 72.9 - 78 50 84 65 79.4 77 74.7 79.1 76.3 80.2 81From above output confidence interval as follows-

s. 95% confidence interval is, I-00, (78.70588 – 75.82353) + t21717-297 * 7.8544 C os , 2.882 35 + 1.69389 7. 85444 0.3429) (